### v1

@@ 5.2.3 Open problems
Similarly, the classical capacity of a quantum channel is given by regularizing the Holevo quantityi $\chi$
@@
In Cubitt et al. (2015a), some evidence is given concerning the uncomputability of the quantum capacity. It is shown that, for any given $n$, there exist channels $T$ with positive quantum capacity that, however, fulfill
@@ 5.3 Fluid mechanics
Motivated by his breakthrough paper (Tao (2016)), Terence Tao started a program (Tao (2017, 2018, 2019a, 2019b)) to address the regularity problem for Navier-Stokes equations in fluid mechanics (one of the Clay’s millennium problems) by showing that fluid mechanics is Turing complete. This would allow encoding a self-replicating machine that blows up in finite time, into the evolution of a fluid. Indeed, Tao (2016) showed this to be possible for an artificial modification (an averaged version) of the Navier-Stokes equations.
A major achievement of that program appeared in a work by Cardona et al. (2021), where they managed to show that there exist Turing complete fluid flows on a Riemannian 3-dimensional sphere. This is done however in the simpler case of Euler equations, which model inviscid incompressible fluids, as opposed to the viscid case treated by the Navier-Stokes equations.
The particular result proven by Cardona et al. (2021) is the existence of fluid flows $X$ in the $3$-dimensional sphere $S_{3}$ so that the following problem is equivalent to the halting problem of a Turing Machine: Given a point $p\in S_{3}$ and an open set $U\subset S_{3}$, decide if the orbit of $X$ through $p$ intersects $U$.
The proof uses deep results of contact topology, together with ideas that were already developed by Moore (1990, 1991) in his construction of uncomputable dynamical systems.
In Cardona et al. (2023), the authors have managed to extend the validity of their result to the Euclidean space $\mathbb{R}^{3}$
@@ 5.4 Quantum Field Theory
We already commented in Chapter 4 on the pioneering work of Komar (1964) about undecidability in quantum field theories. Using a different approach Tachikawa (2023) proved that it is undecidable whether a given 2D supersymmetric Lagrangian theory breaks supersymmetry or not by making a connection with the undecidability of Diophantine equations (section 3.4.1).
@@ 6.1 General view
    -DS1. Fredkin and Toffoli (1982).
    -DS2. Moore (1990).
    -DS3. Omohundro (1984).
    -DS4. da Costa and Doria (1991a, b).
@@
    -SS1. Kanter (1990), section 4.
    -SS2. Gu et al. (2009); Gu and Perales (2012), section 5.1.1.
@@
    -QI1. Brukner (2009); Paterek et al. (2010); Trejo et al. (2023).
    -QI2. Eisert et al. (2012).
    -QI3. Slofstra (2019, 2020).
    -QI4. Fritz (2021).
    -QI5. Mančinska and Roberson (2020).
    -QI6. Fu et al. (2021).
    -QI7. Elkouss and Pérez-García (2018).
    -QI8. Ji et al. (2020).
    -QI9. Wright and Farkas (2023).
    -QI10. Bendersky et al. (2017).
    -QI11. Van den Nest and Briegel (2008).
@@
    -MB1. Cubitt et al. (2015a, 2022); Bausch et al. (2021, 2020).
    -MB2. Watson et al. (2022).
    -MB3. Shiraishi and Matsumoto (2021, 2020).
    -MB4. Purcell et al. (2024).
@@
    -TN1. Morton and Biamonte (2012).
    -TN2. Scarpa et al. (2020).
    -TN3. Kliesch et al. (2014).
    -TN4. De las Cuevas et al. (2016).
@@
    -QF1. Komar (1964), section 4.
    -QF2. Tachikawa (2023), section 5.4.
@@
    -QG1. Geroch and Hartle (1986).
Let’s briefly discuss some conclusions that we can draw from Fig. 1. First, we can appreciate the great diversity of physical systems in which undecidable properties have been proven, both quantum and classical, with most of these results obtained in the last 15 years. We also note that most of them introduce undecidability through mathematical systems derived directly or indirectly from the halting problem for Turing Machines, as in DS1-2, FD1, QI7-10 and MB1-3. As mentioned above, one reason could be that Turing Machines already resemble real physical systems, whereas Gödel’s or AIT arguments are of more abstract nature.
Using the halting problem as starting point, undecidability can be transferred to other mathematical systems by embedding a Turing Machine into them. In this way, one obtains undecidable properties in 2D mathematical systems like cellular automata and tilings that geometrically resemble 2D physical systems (classical or quantum spin systems, tensor networks), so it is possible to export, directly or indirectly, undecidable problems, as in DS3, SS1, SS2, MB1-2, TN1-2, and Q11.
In addition to geometric similarity, the frontier between undecidability in mathematics and physics is crossed by means of other mathematical objects that appear naturally in physics, such as equations, matrices or groups. This is the case for DS4, QI2, QI3-6, TN3-4, QC1, QG1, QF2, which are based on Diophantine Equations, Matrix Mortality or the Word Problem.
@@ 6.2 Infinite undecidable mathematics vs finite decidable nature
Physical systems in nature are effectively finite. Mathematical models that treat them as such are therefore decidable, since it suffices to consider one by one all possible configurations. Take, for instance, 10 particles that are described by an effectively finite configuration space and that interact a finite number of times. Clearly, there is no way to embed a universal Turing Machine with its infinite tape into such a system.
@@
In modelling nature we take limits in other directions as well, which may be displayed by infinitesimal calculus, or asymptotic times—and again, these can either simply the mathematical model or lead to undecidability.
@@
    The associated problem for finite size or finite precision becomes hard/complete for some computational complexity class. Which complexity class will depend on the problem. For instance, the finite-size version of the problem to decide whether a Matrix Product Operator is semidefinite positive is proven to be NP-hard in Kliesch et al. (2014), while the problem of giving an estimate for the ground state energy density of a fixed quantum system in the thermodynamic limit, as a function of the number of bits of precision, is proven to be hard for the complexity class ${\rm FEXP}^{\rm NEXP}$ Aharonov and Irani (2022); Watson and Cubitt (2022). It would be desirable to have general results that allow to infer finite size/precision complexity results based on the undecidability of its infinite counterpart. A first step in that direction is made in Klingler et al. (2023) and discussed further in Wolf (2023).
@@
    As we will discuss in more detail in section 6.5, undecidability of infinite problems may imply unexpected and exotic size-depdendent effects for their finite-size counterparts. That is, there may be a finite but uncomputably large size at which the properties of the system change abruptly and dramatically.
@@ 6.3 The effect of noise
There is yet another important ingredient, which is present in real physical systems and which is usually absent from undecidability results: noise. For instance, it has been shown by Dong et al. (2023), culminating a series of previous works by Qin and Yao (2021, 2023), that, in the presence of noise, the undecidability of approximating the quantum value of a non-local game downgrades to (containment in) the complexity class NEXP. It is an important open question which undecidable results, if any, are stable under natural sources of noise in the corresponding physical problem; and, in case they are not stable, to determine the exact complexity class to which their noisy versions belong.
@@ 6.4 Undecidability over the real numbers
However, in the seminal book by Blum et al. (1998), motivated by optimization algorithms like Newton’s algorithm, the paradigm of computation over the real numbers (instead of the integers) is formalized. In this paradigm, decidable sets are tightly connected with algebraic sets (i.e. zeroes of polynomials). In particular, this allows to show that membership in fractal sets like the Mandelbrot set are undecidable in this different computational paradigm.
The existence of fractal sets in many interesting physical systems, such as Hofstadter’s Butterfly (discovered by Hofstadter (1976)), makes it tempting to suggest that (un)decidability over the reals could also be of relevance in physics.
@@ 6.5 The positive side
Undecidability puts limits on our ability to reason mathematically about a system. But it also has a positive side. It shows the existence of “pathological” examples that can provide counterexamples to open problems in physics. That is, undecidability can be used as a proof technique. We have already seen a couple of examples in this spirit. In section 5.1.3, this technique is used to show that there exist Matrix Product Density Operators without a local purification, a question that goes back to Fannes et al. (1992). In a similar vein, undecidability of the ground state energy density has been used as a proof element in Blakaj and Wolf (2024) in a discussion of the set of reduced states of infinite quantum spin chains. But it is in section 5.2.2 where this technique shows its real power, providing the solution to Tsirelson’s problem and, even more remarkably, to Connes’ embedding problem.
Moreover, those pathological examples could be interesting on their own, since they can display new physics. One illustrative example in this direction is the existence of size-driven quantum phase transitions of Bausch et al. (2018), a new type of quantum effect that is a consequence of the undecidability of the spectral gap: systems whose thermodynamic properties in the infinite limit differ radically from the properties they display at any finite size below a threshold that can be in principle as large as desired. Finding analogous size-driven results for other undecidable problems could be very interesting. This could include for instance be a communication channel whose properties change abruptly after a fixed (but arbitrary large) number of uses. (A step in this direction is shown in Cubitt et al. (2015a), but there the result concerns a family of channels of increasing dimension rather than a single, fixed communication channel.) Or a non-local game for which the quantum value enjoys a fixed jump if played with quantum entanglement dimension larger than a fixed (but arbitrary large, possibly uncomputable) number.

### v2

@@ 5.2.3 Open problems
Similarly, the classical capacity of a quantum channel is given by regularizing the Holevo quantity $\chi$
@@
In Cubitt et al. (2015) [175], some evidence is given concerning the uncomputability of the quantum capacity. It is shown that, for any given $n$, there exist channels $T$ with positive quantum capacity that, however, fulfill
@@ 5.3 Fluid mechanics
Tao (2016) [176] started a program to address the regularity problem for Navier-Stokes equations in fluid mechanics (one of the Clay’s millennium problems) by showing that fluid mechanics is Turing complete [177, 178, 179, 180]. This would allow encoding a self-replicating machine that blows up in finite time, into the evolution of a fluid. Indeed, Tao showed this to be possible for an artificial modification (an averaged version) of the Navier-Stokes equations [176].
A major achievement of that program appeared in a work by Cardona et al. (2021) [181], where they managed to show that there exist Turing complete fluid flows on a Riemannian 3-dimensional sphere. This is done however in the simpler case of Euler equations, which model inviscid incompressible fluids, as opposed to the viscid case treated by the Navier-Stokes equations.
The particular result proven in [181] is the existence of fluid flows $X$ in the $3$-dimensional sphere $S_{3}$ so that the following problem is equivalent to the halting problem of a Turing Machine: Given a point $p\in S_{3}$ and an open set $U\subset S_{3}$, decide if the orbit of $X$ through $p$ intersects $U$.
The proof uses deep results of contact topology, together with the results of Moore [107, 108] commented in Chapter 4, in his construction of uncomputable dynamical systems.
Cardona et al. (2023) [182] have managed to extend the validity of their result to the Euclidean space $\mathbb{R}^{3}$.
@@ 5.4 Quantum Field Theory
We already commented in Chapter 4 on the pioneering work of Komar [3] about undecidability in quantum field theories. Using a different approach Tachikawa (2023) [64] proved that it is undecidable whether a given 2D supersymmetric Lagrangian theory breaks supersymmetry or not by making a connection with the undecidability of Diophantine equations (section 3.4.1).
@@ 6.1 General view
    -DS1. Fredkin and Toffoli (1982) [105].
    -DS2. Moore (1990) [107].
    -DS3. Omohundro (1984) [95].
    -DS4. da Costa and Doria (1991) [62, 25].
    -DS5. Blondel (2001) [114, 115]
@@
    -SS1. Kanter (1990) [81], section 4.
    -SS2. Gu et al. (2009); Gu and Perales (2012) [96, 97], section 5.1.1.
@@
    -QI1. Brukner (2009); Paterek et al. (2010); Trejo et al. (2023) [42, 43, 44].
    -QI2. Eisert et al. (2012) [70].
    -QI3. Slofstra (2019, 2020) [74, 75].
    -QI4. Fritz (2021) [76].
    -QI5. Mancinska and Roberson (2020) [77].
    -QI6. Fu et al. (2021) [78].
    -QI7. Elkouss and Pérez-García (2018 [148].
    -QI8. Ji et al. (2020) [167].
    -QI9. Wright and Farkas (2023) [172].
    -QI10. Bendersky et al. (2017) [173].
    -QI11. Van den Nest and Briegel (2008) [141].
@@
    -MB1. Cubitt et al. (2015, 2022); Bausch et al. (2021, 2020) [175, 120, 127, 126].
    -MB2. Watson et al. (2022) [128].
    -MB3. Shiraishi and Matsumoto (2021, 2020) [131, 132].
    -MB4. Purcell et al. (2024) [45].
@@
    -TN1. Morton and Biamonte (2012) [73].
    -TN2. Scarpa et al. (2020 [83].
    -TN3. Kliesch et al. (2014) [68].
    -TN4. De las Cuevas et al. (2016) [69].
@@
    -QF1. Komar (1964) [3], section 4.
    -QF2. Tachikawa (2023) [64], section 5.4.
@@
    -QG1. Geroch and Hartle (1986) [4].
Let us briefly discuss some conclusions that we can draw from Fig. 1. First, we can appreciate the great diversity of physical systems in which undecidable properties have been proven, both quantum and classical, with most of these results obtained in the last 15 years. We also note that most of them introduce undecidability through mathematical systems derived directly or indirectly from the halting problem for Turing Machines, as in DS1,2,5, FD1, QI7-10 and MB1-3. As mentioned above, one reason is that Turing Machines already resemble real physical systems, whereas Gödel’s or AIT arguments are of more abstract nature.
Using the halting problem as starting point, undecidability can be transferred to other mathematical systems by embedding a Turing Machine into them. In this way, one obtains undecidable properties in 2D mathematical systems like cellular automata and tilings that geometrically resemble 2D physical systems (classical or quantum spin systems, tensor networks), so it is possible to export, directly or indirectly, undecidable problems, as in DS3, SS1, SS2, MB1-2, TN2, and Q11.
In addition to geometric similarity, the frontier between undecidability in mathematics and physics is crossed by means of other mathematical objects that appear naturally in physics, such as equations, matrices or groups. This is the case for DS4, QI2, QI3-6, TN1,3,4, QC1, QG1, QF2, which are based on Diophantine Equations, Matrix Mortality or the Word Problem.
@@ 6.2 Infinite undecidable mathematics vs finite decidable nature
When studying a real physical system, by constraints in the measurement precision or in the sample size, one usually needs to consider only a finite number of possible configurations. Mathematical models to treat those situations as such are therefore decidable, since it suffices to consider one by one all the possible configurations. Take, for instance, 10 particles that are described by an effectively finite configuration space and that interact a finite number of times. Clearly, there is no way to embed a universal Turing Machine with its infinite tape into such a system.
@@
In modeling nature we take limits in other directions as well, which may be displayed by infinitesimal calculus, or asymptotic times—and again, these can either simplify the mathematical model or lead to undecidability.
@@
    The associated problem for finite size or finite precision becomes complete for some computational complexity class. Which complexity class will depend on the problem. For instance, the finite-size version of the problem to decide whether a Matrix Product Operator is semidefinite positive was proven to be NP-hard by Kliesch et al. (2014) [68], while the problem of giving an estimate for the ground state energy density of a fixed quantum system in the thermodynamic limit, as a function of the number of bits of precision, was proven to be hard for the complexity class ${\rm FEXP}^{\rm NEXP}$ (Aharonov and Irani (2022); Watson and Cubitt (2022) [183, 184]). It would be desirable to have general results that allow to infer finite size/precision complexity results based on the undecidability of its infinite counterpart. A first step in that direction was made by Klingler et al. (2023) [185] and discussed further by Wolf (2023) [186].
@@
    As we will discuss in more detail in section 6.5, undecidability of infinite problems may imply unexpected and exotic size-dependent effects for their finite-size counterparts. That is, there may be a finite but uncomputably large size at which the properties of the system change abruptly and dramatically.
@@ 6.3 The effect of noise
There is yet another important ingredient to be considered, which is present in real physical systems: noise. For instance, it has been shown by Dong et al. (2023) [187], culminating a series of previous works by Qin and Yao (2021, 2023) [188, 189], that, under a particular type of noise, where the players are allowed to share arbitrarily many copies of a single state with a fixed amount of noise per state, the undecidability of approximating the quantum value of a non-local game downgrades to (containment in) the complexity class NEXP. It is an important open question to understand which undecidability results are stable under natural sources of noise in the corresponding physical problem; and, in case they are not stable, to determine the exact complexity class to which their noisy versions belong.
@@ 6.4 Undecidability over the real numbers
However, in the seminal book by Blum et al. (1998) [190], motivated by optimization algorithms like Newton’s algorithm, the paradigm of computation over the real numbers (instead of the integers) is formalized. In this paradigm, decidable sets are tightly connected with algebraic sets (i.e. zeroes of polynomials). In particular, this allows to show that membership in fractal sets like the Mandelbrot set are undecidable in this different computational paradigm.
The existence of fractal sets in many interesting physical systems, such as Hofstadter’s Butterfly [191], makes it tempting to suggest that undecidability over the reals could also be of relevance in physics.
@@ 6.5 The positive side
Undecidability puts limits on our ability to reason mathematically about a system. But it also has a positive side. It shows the existence of “pathological” examples that can provide counterexamples to open problems in physics. That is, undecidability can be used as a proof technique. We have already seen a couple of examples in this spirit. In section 5.1.3, this technique is used to show that there exist Matrix Product Density Operators without a local purification, a question that goes back to [192]. In a similar vein, undecidability of the ground state energy density has been used as a proof element by Blakaj and Wolf (2024) [193] in a discussion of the set of reduced states of infinite quantum spin chains. But it is in section 5.2.2 where this technique shows its real power, providing the solution to Tsirelson’s problem and, even more remarkably, to Connes’ embedding problem.
Moreover, those pathological examples could be interesting on their own, since they can display new physics. One illustrative example in this direction is the existence of size-driven quantum phase transitions by Bausch et al. (2018) [194], a new type of quantum effect that is a consequence of the undecidability of the spectral gap: systems whose thermodynamic properties in the infinite limit differ radically from the properties they display at any finite size below a threshold that can be in principle as large as desired. Finding analogous size-driven results for other undecidable problems could be very interesting. This could include for instance a communication channel whose properties change abruptly after a fixed (but arbitrary large) number of uses. (A step in this direction is shown in [175], but there the result concerns a family of channels of increasing dimension rather than a single, fixed communication channel.) Or a non-local game for which the quantum value enjoys a fixed jump if played with quantum entanglement dimension larger than a fixed (but arbitrary large, possibly uncomputable) number.